OpenStudy (anonymous):

Given L_1=<-2,3,-1>+t<2,-3,-5> and 2x1-3x2-5x3=5 Find the intersection point between the line and the plane.

OpenStudy (experimentx):

x1 = -2+2t <-- put these values in the plane and find the value of t.

OpenStudy (anonymous):

What should i do with x2 and x3?

OpenStudy (anonymous):

x2=3-3t?

OpenStudy (anonymous):

and x3=-1-5t?

OpenStudy (experimentx):

yeah!!

OpenStudy (anonymous):

ok one second...

OpenStudy (anonymous):

Okay, I got t=13/38. So plugging that back into my line equation, I get the point <-25/19,75/38,65/38>.

OpenStudy (experimentx):

well ... put those point on the plane and check if it's correct or not.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

@experimentX Could you do work out this problem and do a quick check of the values? I followed your instructions but when I plug them back in to the line equation and the plane equation, I don't get the same position vector....

OpenStudy (experimentx):

sorry ... had connection problem.

OpenStudy (anonymous):

that's okay, solve my problem, like just now lol

OpenStudy (anonymous):

solved*

OpenStudy (anonymous):

i made a sign error....darn. Took forever to find it lol

OpenStudy (experimentx):

oh ... good for you!!